# Spin mechanism of drag resistance in strongly-correlated electron liquids

### v1

@@ I Introduction and motivation
Transport measurements of Coulomb drag provide a useful probe of correlation effects in electron bilayers (see review Narozhny and Levchenko (2016) and references therein). For example, drag resistance $\rho_{\text{D}}$ reveals signatures of superconducting fluctuations Giordano and Monnier (1994), interlayer exciton condensation Seamons et al. (2009); Morath et al. (2009), and exquisite properties of quantum Hall liquids Kellogg et al. (2002); Tutuc et al. (2009). In the present paper, we focus on the physics of nonlocal transport in 2DES at large $r_{s}$, motivated in part by a long-standing problem: the unusual temperature ($T$) and magnetic field ($H$) dependences of drag resistance $\rho_{\text{D}}(T,H)$ observed in bilayer $p$-type GaAs heterojunction devices Pillarisetty et al. (2002, 2003, 2005). For the problem in question, the field dependence concerns the spin effect rather than the orbital effect. We describe a new mechanism of spin-induced magnetodrag resistance, derive the corresponding results that are nonperturbative in interaction, and, in principle, applicable to non-Fermi liquids.
This paper is structured as follows. In Sec. II, we begin with a brief summary of the experimental situation and formulate the problem. We also review the relevant existing theories devoted to this particular topic in parallel. This enables us to place our work in the context of existing studies. In Sec. III, we formulate our approach to the problem of drag in 2DES with $r_{s}\gg 1$. It is based on the theory of hydrodynamic fluctuations. To describe magnetotransport properties, we generalize this theory to include fluctuations of the spin density and spin currents in the presence of partial spin polarization induced by the external field. In Sec. IV, we apply this theory to the calculation of drag magnetoresistance. Finally, in Sec. V, we provide summary of main findings and concluding remarks.
@@ II.1 Temperature dependence
An alternative explanation for the strong enhancement of drag resistance at large $r_{s}$ was proposed in Ref. Spivak and Kivelson (2005), based on the picture of microemulsion phases. At the level of mean field theory, these phases can be characterized as a mixture of microphase-separated regions of liquid and crystal. Practical considerations for the drag resistance were given for the bubble phase, which can be viewed as a suspension of WC islands floating in a uniform Fermi liquid. Under such assumptions, an increased drag can be attributed to the Pomeranchuk effect: because the WC has a higher spin entropy density than a Fermi liquid, the areal fraction of the WC increases with temperature. This fraction can be determined from thermodynamic considerations by examining the free energy difference between the uniform WC and FL phases, which scales linearly with temperature as $\propto nT\ln 2$. This explanation is plausible, although its underlying theory is purely phenomenological. For completeness, we should also mention Ref. Braude and Stern (2001), where an asymmetric WC-FL bilayer was considered with the WC phase pinned by disorder. In this case, drag resistance was shown to scale as $\rho_{\text{D}}\propto T^{4}$. This result is analogous to the Bloch law for the contribution of electron-phonon interaction to the resistivity and drag Bønsager et al. (1998). However, this specific mechanism leads to a significant decrease in transresistivity and requires disorder, thus unlikely to be relevant in application to experiments in high-mobility systems.
@@
It should be emphasized, that even in a weakly-correlated regime, $r_{s}\sim 1$, where the framework of FL applies, the temperature dependence of drag resistance is not simply quadratic. It turns out, Eq. (1) is valid only at lowest temperatures when $T<E_{\text{F}}/(k_{\text{F}}d)$. Note that for GaAs quantum wells typically $k_{\text{F}}d>1$. At slightly higher temperatures, $E_{\text{F}}/(k_{\text{F}}d)<T<E_{\text{F}}/\sqrt{k_{\text{F}}d}$, drag resistance is given by a different expression such that $\rho_{\text{D}}\propto T$. It was implicit in the analysis of Ref. Jauho and Smith (1993), but not clearly emphasized, and later rediscovered in Ref. Chen et al. (2015). As compared to Eq. (1), in this limit, the kinematics of collisions near the threshold of particle-hole continuum for typical frequencies and wave-numbers where $\omega\sim v_{\text{F}}q$ changes the drag resistance to be linear in temperature. The upper limit in the applicability condition of the $\rho_{\text{D}}\propto T$ behavior, namely $\sim E_{\text{F}}/\sqrt{k_{\text{F}}d}$, marks the energy scale at which the intralayer mean free path due to electron collisions $\ell(T)$ becomes comparable or smaller than the interlayer distance $d$. This regime defines the onset of collision-dominated transport. At even higher temperatures, $E_{\text{F}}/\sqrt{k_{\text{F}}d}<T<E_{\text{F}}/\sqrt[4]{k_{\text{F}}d}$, drag resistance is dominated by plasmons and $\rho_{\text{D}}\propto T^{3}$. The energy scale $\sim E_{\text{F}}/\sqrt[4]{k_{\text{F}}d}$ corresponds to the hydrodynamic regime of plasmons, when their frequency $\omega_{\text{pl}}$ at the typical wave-numbers, $q\sim 1/d$, becomes comparable or smaller than the rate of electron collisions $\tau^{-1}$. Above this temperature, but still below $E_{\text{F}}$, one finds drag decreasing with temperature as $\rho_{D}\propto 1/T$ Apostolov et al. (2014).
@@ II.4 Hydrodynamic description
In this work, we develop a theory of drag in bilayers of spin-polarized electron liquids. Our approach is based on the hydrodynamic description of transport in the regime of strong electron correlations originally described in Ref. Andreev et al. (2011). This approach can be justified in high-mobility conductors within a range of temperatures and sample purities where the electron-electron mean-free path is small compared to the length scales over which momentum and energy conservation of electron systems are violated. An extension of the theory in Ref. Andreev et al. (2011) to include hydrodynamic fluctuations and compute drag resistance in zero field was given in Refs. Apostolov et al. (2014); Chen et al. (2015); Holder (2019). The orbital effect of the magnetic field was considered in Refs. Patel et al. (2017); Apostolov et al. (2019). Here, we focus on the spin effect and incorporate fluctuations of the spin density and spin current into the main set of hydrodynamic equations, supplemented by the fluctuation-dissipation relations. This enables us to consider both drag and single-layer resistivities from a unified perspective. Our results are as general as any hydrodynamic theory and rely only on the applicability conditions of the hydrodynamic approximation. We analyze our results for the temperature, field, and density dependence of drag resistance in light of experimental findings Pillarisetty et al. (2002, 2003, 2005).
@@ III.1 Main equations
It is important to stress that only the spin component along the magnetic field appears in this equation since the other spin components are not conserved due to spin precession. The fluctuation of the spin current density
@@
is related to the gradient of the spin chemical potential fluctuation $\delta\mu_{\sigma}$ via the spin conductivity $\sigma_{s}$. The latter can be expressed in terms of the spin diffusion constant $D_{\sigma}$ via the Einstein relation $\sigma_{s}=\chi D_{\sigma}$, where $\chi$ is the magnetic susceptibility. In the expression for the spin current, we neglected the term corresponding to the spin thermocurrent that must be present by the Onsager principle. Its contribution is important for spin-caloric and spin-drag thermal resistances near charge neutrality Levchenko and Andreev (2024); Levchenko (2024), but is expected to be small for electrical properties at high-doping as compared to the spin diffusion term. In accordance with the fluctuation-dissipation theorem, the random flux of Langevin spin currents correlates to the spin conductivity
@@
The hydrodynamic description applies to any liquid type whether it is Fermi liquid or not. Microscopic properties of the liquid manifest themselves via the temperature and density dependence of the intrinsic dissipative coefficients as well as thermodynamic quantities, and the equation of state. Noteworthy, in high-mobility and low-density 2DES the momentum and energy relaxing electron-phonon scattering is still weak even at $T>E_{\text{F}}$ Gao et al. (2005). Therefore, the hydrodynamic description applies from very short distances of the order of inter-electron spacing for both semiquantum liquids at $E_{\text{F}}<T<\sqrt{r_{s}}E_{\text{F}}$, or highly correlated classical fluid at $\sqrt{r_{s}}E_{\text{F}}<T<r_{s}E_{\text{F}}$. It should be mentioned that there is no established theory in these regimes, but a conjecture concerning the $T$ dependences of $s,\kappa,\eta,\zeta$ was put forward in Ref. Spivak and Kivelson (2006). We have also our own results for some of these quantities that will be reported in a separate publication.
@@ III.2 Solution for a bilayer setup
Consider now a bilayer geometry relevant for the Coulomb drag setup Fig. 1. For simplicity we assume identical 2DES. To describe this system we need to duplicate the number of equations describing fluctuations of particle density Eq. (2), entropy density Eq. (4), spin density Eq. (7), and hydrodynamic velocity Eq. (10) in each of the layer and their coupling. For hydrodynamic model to be applicable, we are constrained to work under the tacit condition that interlayer separation exceeds the intralayer mean free path due to electron collisions $d\gg\ell$. Since $\ell$ diverges in the low temperature limit, hydrodynamics sets in at the intermediate range of temperatures. For a Fermi liquid, the onset of collision-dominated regime is marked by the temperature $T\gtrsim E_{\text{F}}/\sqrt{k_{\text{F}}d}$. In the following, we denote density fluctuations in each layer as $\delta n_{1,2}(\bm{r},t)$, and similarly for all other quantities.
@@
$[-i\omega+i(\bm{q}\cdot\bm{v})]\delta\sigma_{1}=-\sigma_{\sigma}q^{2}\delta\mu_{\sigma 1}-i(\bm{q}\cdot\delta\bm{\varsigma}_{1}).$ (16)
@@
To close the system of required relations for the obtained solution, we need expressions for thermal averages, which follow directly from Eqs. (9) and (III.1), namely
@@ IV Spin drag magnetoresistance
Figure 2: Dispersion laws $\omega=\omega_{\pm}(q)$ for the optical and acoustic plasmons Eq. (25), shown by dashed lines, superimposed on top of the color plot that defines magnitude of the integrand in Eq. (30), which was normalized to itself at the maximal value. The lower dashed line defines broadening of the spectral weight due to spin diffusion. This plot was generated for values $\alpha=1.5\times 10^{-3}$ and $\gamma=2.5\times 10^{-2}$.
@@
$\rho_{\sigma}\simeq\frac{\sigma_{s}}{e^{2}}\left(\frac{\epsilon v_{F}}{e^{2}}\right)^{2}\frac{f(\gamma)}{(nd^{2})^{3}}\frac{T}{E_{\text{F}}}\left(\frac{H}{E_{\text{F}}}\right)^{2},$ (33)
@@
The field independent part of drag resistance, can be calculated in the same fashion. Indeed, using thermal averages
@@
where $\Lambda_{q}=4T(\eta+\zeta)q^{4}/m^{2}$. This leads then to a corresponding part of the density structure factor computed to the linear order in $\bm{v}$
@@
Drag resistance defined by Eqs. (33) and (38) are the main results of this work.
@@ V Summary and discussion
$\bullet$ The measured spin diffusion constant via spin Coulomb drag in 2DES Weber et al. (2005) reveals that $D_{\sigma}$ is almost temperature independent in a broad range $T>E_{\text{F}}/10$. With this input, Eq. (33) predicts that drag magnetoresistance is linear in temperature, $\rho_{\sigma}\propto T$. This temperature behavior is qualitatively consistent with the observations Pillarisetty et al. (2003, 2005).
$\bullet$ The plasmon enhancement factor described by $f(\gamma)$ is also temperature dependent, but only logarithmically. Indeed, it can be readily verified that for $T<E_{\text{F}}$ the parameter $\gamma<1$. The asymptote of Eq. (34) in the limit $\gamma\ll 1$ is $f\sim 2\ln^{6}[\gamma\ln(1/\gamma)]$. For a Fermi liquid $\eta\sim n(E_{\text{F}}/T)^{2}$, therefore $f$ behaves logarithmically with $T$. Note, however, that numerically $f\sim 25$ for $\gamma\sim 0.1$ so that plasmon enhancement is significant and strongest at matched density between the layers. The corresponding enhancement for the field independent part, as described by $h(\gamma)$ function in Eq.(39), is slightly weaker, see Fig. 3 for comparison.
$\bullet$ Eq. (33) also predicts quadratic field dependence, $\rho_{\sigma}\propto H^{2}$, which we expect to saturate at higher fields since spin polarization saturates. Additional complications may arise from the $H$ dependence of the kinetic coefficients $\eta(H),\zeta(H)$.
$\bullet$ As a function of density, the relevant prefactor in Eq. (33) scales as $\rho_{\sigma}\propto\sigma_{s}(n)/n^{5}$. Provided that spin conductivity does not change significantly in the measured range of densities, the power law scaling of drag resistance with $n$ is therefore also in good agreement with the observed dependence.
$\bullet$ The comparison of Eq. (38) to the FL formula for drag resistivity Eq. (1) reveals that intralayer collisions strongly enhance drag effect. This is qualitatively consistent with the fact that drag resistance is bilayers with large $r_{s}$ is found to be two to three orders of magnitude larger than expected on the basis of FL theory Pillarisetty et al. (2002).
$\bullet$ The quantum Monte Carlo results for the spin susceptibility $\chi$ reported Attaccalite et al. (2002) have been widely compared with experiment e.g. Zhu et al. (2003), and appear quite reliable and accurate. These studies show that spin susceptibility of a 2DES, measured in the units of Pauli susceptibility, $\chi/\chi_{\text{P}}$, growths by an order of magnitude when $r_{s}$ is increased. As a consequence, based on the Einstein relation $\sigma_{s}=\chi D_{\sigma}$, Eq. (33) predicts strongly enhanced drag magnetoresistance.
$\bullet$ We note that our results can be extrapolated to the nondegenerate limit $T>E_{\text{F}}$ that includes both semiquantum and strongly correlated classical plasma. For example, for a classical limit $\eta\sim\zeta\sim mv_{T}T/e^{2}$, where $v_{T}\sim\sqrt{T/m}$ is the thermal velocity, so that $\rho_{\nu}\propto T^{5/2}$.

### v2

# Spin mechanism of drag resistance in strongly correlated electron liquids

@@ I Introduction and motivation
Transport measurements of Coulomb drag provide a useful probe of correlation effects in electron bilayers (see review Narozhny and Levchenko (2016) and references therein). For example, drag resistance $\rho_{\text{D}}$ reveals signatures of superconducting fluctuations Giordano and Monnier (1994), interlayer exciton condensation Seamons et al. (2009); Morath et al. (2009), and exquisite properties of quantum Hall liquids Kellogg et al. (2002); Tutuc et al. (2009). In the present paper, we focus on the physics of nonlocal transport in 2DES at large $r_{s}$, motivated in part by a long-standing problem: the unusual temperature ($T$) and magnetic field ($H$) dependences of drag resistance $\rho_{\text{D}}(T,H)$ observed in bilayer $p$-type GaAs heterojunction devices Pillarisetty et al. (2002, 2003, 2005). For the problem in question, the field dependence concerns the spin effect rather than the orbital effect. We describe a particular mechanism of spin-induced magnetodrag resistance, derive the corresponding results that are nonperturbative in interaction, and, in principle, applicable to non-Fermi liquids.
This paper is structured as follows. In Sec. II, we begin with a brief summary of the experimental situation and formulate the problem. We also review the relevant existing theories devoted to this particular topic in parallel. This enables us to place our paper in the context of existing studies. In Sec. III, we formulate our approach to the problem of drag in 2DES with $r_{s}\gg 1$. It is based on the theory of hydrodynamic fluctuations. To describe magnetotransport properties, we generalize this theory to include fluctuations of the spin density and spin currents in the presence of partial spin polarization induced by the external field. In Sec. IV, we apply this theory to the calculation of drag magnetoresistance. Finally, in Sec. V, we provide summary of main findings and concluding remarks.
@@ II.1 Temperature dependence
An alternative explanation for the strong enhancement of drag resistance at large $r_{s}$ was proposed in Ref. Spivak and Kivelson (2005), based on the picture of microemulsion phases. At the level of mean-field theory, these phases can be characterized as a mixture of microphase-separated regions of liquid and crystal. Practical considerations for the drag resistance were given for the bubble phase, which can be viewed as a suspension of WC islands floating in a uniform Fermi liquid. Under such assumptions, an increased drag can be attributed to the Pomeranchuk effect: because the WC has a higher spin entropy density than a Fermi liquid, the areal fraction of the WC increases with temperature. This fraction can be determined from thermodynamic considerations by examining the free energy difference between the uniform WC and FL phases, which scales linearly with temperature as $\propto nT\ln 2$. This explanation is plausible, although its underlying theory is purely phenomenological. For completeness, we should also mention Ref. Braude and Stern (2001), where an asymmetric WC-FL bilayer was considered with the WC phase pinned by disorder. In this case, drag resistance was shown to scale as $\rho_{\text{D}}\propto T^{4}$. This result is analogous to the Bloch law for the contribution of electron-phonon interaction to the resistivity and drag Bønsager et al. (1998). However, this specific mechanism leads to a significant decrease in transresistivity and requires disorder, thus unlikely to be relevant in application to experiments in high-mobility systems.
@@
It should be emphasized, that even in a weakly-correlated regime, $r_{s}\sim 1$, where the framework of FL applies, the temperature dependence of drag resistance is not simply quadratic. It turns out that Eq. (1) is valid only at lowest temperatures when $T<E_{\text{F}}/(k_{\text{F}}d)$. Note that for GaAs quantum wells typically $k_{\text{F}}d>1$. At slightly higher temperatures, $E_{\text{F}}/(k_{\text{F}}d)<T<E_{\text{F}}/\sqrt{k_{\text{F}}d}$, drag resistance is given by a different expression such that $\rho_{\text{D}}\propto T$. It was implicit in the analysis of Ref. Jauho and Smith (1993), but not clearly emphasized, and later rediscovered in Ref. Chen et al. (2015). As compared to Eq. (1), in this limit, the kinematics of collisions near the threshold of particle-hole continuum for typical frequencies and wave numbers where $\omega\sim v_{\text{F}}q$ changes the drag resistance to be linear in temperature. The upper limit in the applicability condition of the $\rho_{\text{D}}\propto T$ behavior, namely $\sim E_{\text{F}}/\sqrt{k_{\text{F}}d}$, marks the energy scale at which the intralayer mean-free path because of electron collisions $\ell(T)$ becoming comparable or smaller than the interlayer distance $d$. This regime defines the onset of collision-dominated transport. At even higher temperatures, $E_{\text{F}}/\sqrt{k_{\text{F}}d}<T<E_{\text{F}}/\sqrt[4]{k_{\text{F}}d}$, drag resistance is dominated by plasmons and $\rho_{\text{D}}\propto T^{3}$. The energy scale $\sim E_{\text{F}}/\sqrt[4]{k_{\text{F}}d}$ corresponds to the hydrodynamic regime of plasmons, when their frequency $\omega_{\text{pl}}$ at the typical wave-numbers, $q\sim 1/d$, becomes comparable or smaller than the rate of electron collisions $\tau^{-1}$. Above this temperature, but still below $E_{\text{F}}$, one finds drag decreasing with temperature as $\rho_{D}\propto 1/T$ Apostolov et al. (2014).
@@ II.4 Hydrodynamic description
In this paper, we develop a theory of drag in bilayers of spin-polarized electron liquids. Our approach is based on the hydrodynamic description of transport in the regime of strong electron correlations originally described in Ref. Andreev et al. (2011). This approach can be justified in high-mobility conductors within a range of temperatures and sample purities where the electron-electron mean-free path is small compared to the length scales over which momentum and energy conservation of electron systems are violated. An extension of the theory in Ref. Andreev et al. (2011) to include hydrodynamic fluctuations and compute drag resistance in zero field was given in Refs. Apostolov et al. (2014); Chen et al. (2015); Holder (2019). The orbital effect of the magnetic field was considered in Refs. Patel et al. (2017); Apostolov et al. (2019). Here, we focus on the spin effect and incorporate fluctuations of the spin density and spin current into the main set of hydrodynamic equations, supplemented by the fluctuation-dissipation relations. This enables us to consider both drag and single-layer resistivities from a unified perspective. Our results are as general as any hydrodynamic theory and rely only on the applicability conditions of the hydrodynamic approximation. We analyze our results for the temperature, field, and density dependence of drag resistance in light of experimental findings Pillarisetty et al. (2002, 2003, 2005).
@@ III.1 Main equations
It is important to stress that only the spin component along the magnetic field appears in this equation since the other spin components are not conserved because of spin precession. The fluctuation of the spin current density
@@
is related to the gradient of the spin chemical potential fluctuation $\delta\mu_{\sigma}$ via the spin conductivity $\sigma_{s}$. The latter can be expressed in terms of the spin diffusion constant $D_{\sigma}$ via the Einstein relation $\sigma_{s}=\chi D_{\sigma}$, where $\chi$ is the magnetic susceptibility. In the expression for the spin current, we neglected the term corresponding to the spin thermocurrent that must be present by the Onsager principle. Its contribution is important for spin-caloric and spin-drag thermal resistances near charge neutrality Levchenko and Andreev (2024); Levchenko (2024), but is expected to be small for electrical properties at high doping as compared to the spin diffusion term. In accordance with the fluctuation-dissipation theorem, the random flux of Langevin spin currents correlates to the spin conductivity
@@
The hydrodynamic description applies to any liquid type whether it is Fermi liquid or not. Microscopic properties of the liquid manifest themselves via the temperature and density dependence of the intrinsic dissipative coefficients as well as thermodynamic quantities, and the equation of state. Noteworthy, in high-mobility and low-density 2DES the momentum and energy-relaxing electron-phonon scattering is still weak even at $T>E_{\text{F}}$ Gao et al. (2005). Therefore, the hydrodynamic description applies from very short distances of the order of interelectron spacing for both semiquantum liquids at $E_{\text{F}}<T<\sqrt{r_{s}}E_{\text{F}}$, or highly correlated classical fluid at $\sqrt{r_{s}}E_{\text{F}}<T<r_{s}E_{\text{F}}$. It should be mentioned that there is no established theory in these regimes, but a conjecture concerning the $T$ dependences of $s,\kappa,\eta,\zeta$ was put forward in Ref. Spivak and Kivelson (2006). We have also our own results for some of these quantities that will be reported in a separate publication.
@@ III.2 Solution for a bilayer setup
Consider now a bilayer geometry relevant for the Coulomb drag setup Fig. 1. For simplicity we assume identical 2DES. To describe this system we need to duplicate the number of equations describing fluctuations of particle density Eq. (2), entropy density Eq. (4), spin density Eq. (7), and hydrodynamic velocity Eq. (10) in each of the layer and their coupling. For hydrodynamic model to be applicable, we are constrained to work under the tacit condition that interlayer separation exceeds the intralayer mean-free path because of electron collisions $d\gg\ell$. Since $\ell$ diverges in the low temperature limit, hydrodynamics sets in at the intermediate range of temperatures. For a Fermi liquid, the onset of collision-dominated regime is marked by the temperature $T\gtrsim E_{\text{F}}/\sqrt{k_{\text{F}}d}$. In the following, we denote density fluctuations in each layer as $\delta n_{1,2}(\bm{r},t)$, and similarly for all other quantities.
@@
$[-i\omega+i(\bm{q}\cdot\bm{v})]\delta\sigma_{1}=-\sigma_{s}q^{2}\delta\mu_{\sigma 1}-i(\bm{q}\cdot\delta\bm{\varsigma}_{1}).$ (16)
@@
To close the system of required relations for the obtained solution, we need expressions for thermal averages, which follow directly from Eqs. (9) and (III.1), namely,
@@ IV Spin drag magnetoresistance
Figure 2: Dispersion laws $\omega=\omega_{\pm}(q)$ for the optical and acoustic plasmons Eq. (25), shown by dashed lines, superimposed on top of the color plot that defines magnitude of the integrand in Eq. (30), which was normalized to itself at the maximal value. The lower dashed line defines broadening of the spectral weight because of spin diffusion. This plot was generated for values $\alpha=1.5\times 10^{-3}$ and $\gamma=2.5\times 10^{-2}$.
@@
$\rho_{\sigma}\simeq\frac{\sigma_{s}}{e^{2}}\left(\frac{\epsilon v_{\text{F}}}{e^{2}}\right)^{2}\frac{f(\gamma)}{(nd^{2})^{3}}\frac{T}{E_{\text{F}}}\left(\frac{H}{E_{\text{F}}}\right)^{2},$ (33)
@@
The field-independent part of drag resistance, can be calculated in the same fashion. Indeed, using thermal averages
@@
where $\Lambda_{q}=4T(\eta+\zeta)q^{4}/m^{2}$. This leads then to a corresponding part of the density structure factor computed to the linear order in $\bm{v}$,
@@
Drag resistance defined by Eqs. (33) and (38) are the main results of this paper.
@@ V Summary and discussion
(i) The measured spin diffusion constant via spin Coulomb drag in 2DES Weber et al. (2005) reveals that $D_{\sigma}$ is almost temperature independent in a broad range $T>E_{\text{F}}/10$. With this input, Eq. (33) predicts that drag magnetoresistance is linear in temperature, $\rho_{\sigma}\propto T$. This temperature behavior is qualitatively consistent with the observations Pillarisetty et al. (2003, 2005).
(ii) The plasmon enhancement factor described by $f(\gamma)$ is also temperature dependent, but only logarithmically. Indeed, it can be readily verified that for $T<E_{\text{F}}$ the parameter $\gamma<1$. The asymptote of Eq. (34) in the limit $\gamma\ll 1$ is $f\sim 2\ln^{6}[\gamma\ln(1/\gamma)]$. For a Fermi liquid $\eta\sim n(E_{\text{F}}/T)^{2}$, therefore $f$ behaves logarithmically with $T$. Note, however, that numerically $f\sim 25$ for $\gamma\sim 0.1$ so that plasmon enhancement is significant and strongest at matched density between the layers. The corresponding enhancement for the field independent part, as described by $h(\gamma)$ function in Eq.(39), is slightly weaker, see Fig. 3 for comparison.
(iii) Equation (33) also predicts quadratic field dependence, $\rho_{\sigma}\propto H^{2}$, which we expect to saturate at higher fields since spin polarization saturates. Additional complications may arise from the $H$ dependence of the kinetic coefficients $\eta(H),\zeta(H)$.
(iv) As a function of density, the relevant prefactor in Eq. (33) scales as $\rho_{\sigma}\propto\sigma_{s}(n)/n^{5}$. Provided that spin conductivity does not change significantly in the measured range of densities, the power law scaling of drag resistance with $n$ is therefore also in good agreement with the observed dependence.
(v) The comparison of Eq. (38) to the FL formula for drag resistivity Eq. (1) reveals that intralayer collisions strongly enhance drag effect. This is qualitatively consistent with the fact that drag resistance is bilayers with large $r_{s}$ is found to be two to three orders of magnitude larger than expected on the basis of FL theory Pillarisetty et al. (2002).
(vi) The quantum Monte Carlo results for the spin susceptibility $\chi$ reported Attaccalite et al. (2002) have been widely compared with experiment e.g. Zhu et al. (2003), and appear quite reliable and accurate. These studies show that spin susceptibility of a 2DES, measured in the units of Pauli susceptibility, $\chi/\chi_{\text{P}}$, grows by an order of magnitude when $r_{s}$ is increased. As a consequence, based on the Einstein relation $\sigma_{s}=\chi D_{\sigma}$, Eq. (33) predicts strongly enhanced drag magnetoresistance.
(vii) We note that our results can be extrapolated to the nondegenerate limit $T>E_{\text{F}}$ that includes both semiquantum and strongly correlated classical plasma. For example, for a classical limit $\eta\sim\zeta\sim mv_{T}T/e^{2}$, where $v_{T}\sim\sqrt{T/m}$ is the thermal velocity, so that $\rho_{\nu}\propto T^{5/2}$.
@@
The result for drag magnetoresistance in Eq. (27) can be generalized to systems with broken Galilean invariance. The key modification required is the inclusion of the electrical current because of finite intrinsic conductivity $\sigma_{c}$ (i.e., charge conductivity relative to the fluid, often referred to as microscopic “incoherent” conductivity). Consequently, Langevin fluxes must be augmented to account for fluctuations arising from intrinsic currents in the fluid. These fluctuations introduce an additional mechanism for density variation via current continuity, thereby contributing to the dynamical structure factor of the fluid and the interlayer drag force. While the drag effect is still enhanced by plasmons, their lifetime is now governed by the Maxwell time of charge relaxation rather than viscous diffusion. For long-wavelength density fluctuations, the corresponding attenuation coefficient is given by $\Im\omega_{q}=2\pi\sigma_{c}q/\epsilon$. This change in physics affects the broadening of plasmon resonances and the enhancement factor in drag resistance. The analysis indicates that Eq. (27) acquires an additional contribution of the form
$\rho_{c}=\frac{\sigma_{c}}{4\pi^{2}e^{4}}\frac{T}{E_{\text{F}}}\frac{1}{(nd^{2})^{2}}g(\beta).$ (40)
The dimensionless function $g(\beta)$ captures the plasmon enhancement and exhibits a logarithmic dependence on the dimensionless parameter
$\beta\simeq\frac{\sigma_{c}}{e^{2}}\sqrt{\frac{e^{2}}{\epsilon v_{\text{F}}}}\frac{1}{\sqrt{k_{\text{F}}d}}.$ (41)
The main parametric dependence in Eqs. (33) and (38) remains the same; however, the dimensionless functions $f(\gamma)\to f(\beta)$ and $h(\gamma)\to h(\beta)$ are described by different analytic expressions in terms of $\beta$ as compared to Eqs. (34) and (39) respectively. Therefore, modulo logarithmic factors, the general conclusion regarding the field and temperature dependence of the drag magnetoresistance remains the same in this case as well.